OpenStudy (anonymous):

Trig / Pre Cal/identities Am I on the right path? Please do not give the answer. Here is the problem \( \frac{(\sin x- \cos x)*2}{cos} = \sec x - \cos x\) I am working on the left side and need to prove it equals the right. This is what I have done so far \( \frac{(\sin^2 x - 2(\cos x)(\sin x)+\cos^2 x}{\cos x} = \sec x - \cos x\) Am I going about this the right way because I don't see where to go from here. Thank you.

OpenStudy (anonymous):

It is suppose to be Trig / Pre Cal/identities Am I on the right path? Please do not give the answer. Here is the problem \( \frac{(\sin x- \cos x)^2}{cos} = \sec x - \cos x\) I am working on the left side and need to prove it equals the right. This is what I have done so far \( \frac{(\sin^2 x - 2(\cos x)(\sin x)+\cos^2 x}{\cos x} = \sec x - \cos x\) Am I going about this the right way because I don't see where to go from here. Thank you.

OpenStudy (anonymous):

This is where I am at. Am I on the right path and where do I go from here because I do not see anything. \( \frac{\sin^2 x - 2(\cos x)(\sin x)+\cos^2 x}{\cos x} = \sec x - \cos x\)

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

Made a mistake

Nnesha (nnesha):

hmm well....... i guess both sides are not equal

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nixy Made a mistake \(\color{blue}{\text{End of Quote}}\) what ?? :-)

OpenStudy (anonymous):

It is suppose to be \( \frac{(\sin x- \cos x)^2}{cos} = \sec x - 2\sin x \) And I am at \( \frac{\sin^2 x - 2(\cos x)(\sin x)+\cos^2 x}{\cos x} = \sec x - 2\sin x \)

OpenStudy (anonymous):

I was typing it in too fast and made a boo boo. Sorry about that.

OpenStudy (anonymous):

Try sin² x = 1 - cos² x

OpenStudy (anonymous):

So the numerator becomes \[1 - cos^2 x - 2 (sin x)(cos x) + cos^2 x\]

Nnesha (nnesha):

that's right or just put 1 sin^2x+cos^2 =1 you will get the same answer :-)

OpenStudy (anonymous):

That is it peachpi.

OpenStudy (anonymous):

I did that a min ago and it did not work but re looked at it and I see where I made my mistake. Thank you all

OpenStudy (anonymous):

I think I over think these things at time. One minute I am rolling through them and then I get snagged by the simplest thing lol Thank you all

Nnesha (nnesha):

haha great work!!! :-)